How do I solve this?

The shock absorbers in a car act as a big spring. When a 92.5 kg person gets in, the car sinks 0.0131 m. What is the spring constant of the shocks? (Unit N/m) 003-20Accls Corporation All Rights Reorved
Answer #1

Use the following formula to solve the given problem:

F = kx

=> k = F/x and force F = weight = mg, so

k = mg/x

= (92.5*9.8)/0.0131

= 6.92*10^4 N/m